OpenStudy (anonymous):

how do you solve arctan(x) + x/(x^2+1)?

OpenStudy (anonymous):

I know the answer is x=0 I just don't know how to get there

OpenStudy (mertsj):

There is nothing to solve.

OpenStudy (anonymous):

arctan(x) + x/(x^2+1)=0 is what I meant actually sorry

OpenStudy (anonymous):

I can change it so it looks like x=tan(-x/(x^2+1)), but I'm not sure that helps

OpenStudy (mertsj):

Is this one of those problems where you have to integrate by parts?

OpenStudy (anonymous):

this problem is to find the critical points for x*arctan(x)

OpenStudy (jamesj):

Let \( f(x) = \arctan(x) + x/(x^2 + 1) \). Then \[ f'(x) = \frac{1}{x^2+1} + \frac{1-x^2}{(x^2+1)^2} = \frac{2}{(x^2+1)^2} > 0 \ \ \hbox{for all} \ \ x \] Therefore the function is strictly monotonically increasing. Now f(0) = 0. Given these facts can there be any other \( x \neq 0 \) such that \( f(x) = 0\) ?

OpenStudy (anonymous):

is there any way to use trig to simplify this one?

OpenStudy (jamesj):

I've just given you answer: - f(x) is an everywhere strictly increasing function - f(0) = 0 Given these two facts about f, is it possible there can be another solution? i.e., another x not equal to zero such f(x) = 0?

OpenStudy (mertsj):

What class are you taking? Have you had calculus?

OpenStudy (anonymous):

how can you know that x=0? I am in calculus right now

OpenStudy (anonymous):

other than subbing in 0 and checking

OpenStudy (jamesj):

arctan(0) = 0 and if x = 0, then certainly x/(x^2 + 1) = 0. Therefore f(0) = 0.

OpenStudy (anonymous):

how did you know to check at x=0?

OpenStudy (jamesj):

Because it was obvious value to check. Good luck to you.

OpenStudy (anonymous):

ok, thanks

OpenStudy (mertsj):

So could you post the original problem just as it appears in your textbook>

OpenStudy (anonymous):

I'm looking for the critical points of x*arctan(x)

OpenStudy (anonymous):

so when the derivative is 0 arctan(x) + x/(x^2+1)=0 I know the answer is 0 because I can sub it in and it works, I just wanted to know the way to simplify the probem showing that the derivative is 0 when x=0

OpenStudy (jamesj):

Ok. So for g(x) = x.arctan(x), g'(x) = 0 if and only if x = 0. At that critical value, g(0) = 0. Further, g(x) is an even function, g(x) = g(-x); and for x > 0, g(x) > 0. Therefore g(x) has a minimum at x = 0.

OpenStudy (anonymous):

Ok, I think that is enough justification for x=0, all I wanted to know was if there was a way to simplify arctan(x) + x/(x^2+1)=0 to solve for x

OpenStudy (jamesj):

The answer is there is no way to solve this problem analytically, other than the one I've given. x=0 works, but if you didn't think to guess that then you might be stuck for a while trying to figure it out.

OpenStudy (anonymous):

alright, thanks for your help